OpenStudy (beth12345):

how to solve for feda when cos(theta)=-0.19?

OpenStudy (anonymous):

I'm sorry that wasn't for you.

OpenStudy (anonymous):

I have so many tabs open, I got mixed up.

OpenStudy (anonymous):

And I was just joking around with my freinds.

OpenStudy (beth12345):

ok,np

OpenStudy (anonymous):

theta = anticos (-0.19)

OpenStudy (anonymous):

feda = anticos (-0.19)

OpenStudy (beth12345):

\[\theta=\cos ^{-1}(-0.19)\] like this, and you were actually right the first time, its theda

OpenStudy (beth12345):

I did this but in the back of the book it says the answers are both 1.76 and 4.90 and i dont know how to get it

OpenStudy (beth12345):

The original question was...solve each equation for \[0 \le \theta <2\pi\] Give solutions to the nearest hundredth of a radian

OpenStudy (anonymous):

use a calculator set it to radians